OpenStudy (anonymous):

8. Find the distance between points M(6, 16) and Z(–1, 14) to the nearest tenth.

OpenStudy (anonymous):

7.28

OpenStudy (anonymous):

thnks but how did you get that?

OpenStudy (anonymous):

sqrt [(14 - 16)^2 + (-1+-6 )^2] = sqrt(53) = 7.28

OpenStudy (anonymous):

\[\sqrt{(x2-x1)^2+(y2-y1)^2}\]

OpenStudy (anonymous):

Medal given because you are the first person to ask me how...ever

OpenStudy (anonymous):

basically its pythagoras theorem